# ToMCCA-3: A realistic 3-body coalescence model

### v1

@@ Acknowledgments
$\gamma\frac{\mathrm{d}N_{\mathrm{A}}}{\mathrm{\leavevmode\nobreak\ d}^{3}P}=\frac{S_{\mathrm{A}}}{(2\pi)^{14}}|J^{-1}|^{3}\int\mathrm{d}^{3}q_{1}\int\mathrm{d}^{3}q_{2}\int\mathrm{d}^{3}S\int\mathrm{d}^{3}\xi_{1}\,\int\mathrm{d}^{3}\xi_{2}\int\mathrm{d}^{3}\xi_{1}^{\prime}\int\mathrm{d}^{3}\xi_{2}^{\prime}\\ \times\varphi^{*}_{NR}\left(\bm{\xi}_{1}^{\prime},\bm{\xi}_{2}^{\prime}\right)\varphi_{NR}\left(\bm{\xi}_{1},\bm{\xi}_{2}\right)\,e^{-i{\bm{q}}_{1}\cdot\left(\bm{\xi}_{1}-\bm{\xi}_{1}^{\prime}\right)-i{\bm{q}}_{2}\cdot\left(\bm{\xi}_{2}-\bm{\xi}_{2}^{\prime}\right)}\,\\ \times W_{\mathrm{A}}\left((c_{1}P^{0}_{d},{\bm{p}}_{1}),(c_{2}P^{0}_{d},{\bm{p}}_{2}),(c_{3}P^{0}_{d},{\bm{p}}_{3}),\frac{(t_{0},{\bm{x}}_{1}+{\bm{x}}_{1}^{\prime})}{2},\frac{(t_{0},{\bm{x}}_{2}+{\bm{x}}_{2}^{\prime})}{2},\frac{(t_{0},{\bm{x}}_{3}+{\bm{x}}_{3}^{\prime})}{2}\right)\Bigr{|}_{P=P_{\mathrm{A}}}\,.$ (60)
$t_{0}$ can be taken arbitrarily as $t_{0}=0$. Reinserting the integration over $d^{3}S\times\delta^{(3)}({\bm{R}}-{\bm{R}}^{\prime})$, and going back to the integration over the particle coordinates and momenta, we have
$\gamma\frac{\mathrm{d}N_{\mathrm{A}}}{\mathrm{d}^{3}P}=\frac{S_{\mathrm{A}}}{(2\pi)^{14}}\int\mathrm{d}^{3}p_{1}\int\mathrm{d}^{3}p_{2}\int\mathrm{d}^{3}p_{3}\int\mathrm{d}^{3}x_{1}\int\mathrm{d}^{3}x_{2}\,\int\mathrm{d}^{3}x_{3}\int\mathrm{d}^{3}x_{1}^{\prime}\int\mathrm{d}^{3}x_{2}^{\prime}\int\mathrm{d}^{3}x_{3}^{\prime}\\ \times\Psi_{NR}^{*}\left({\bm{x}}_{1}^{\prime},{\bm{x}}_{2}^{\prime},{\bm{x}}_{3}^{\prime},\right)\Psi_{NR}\left({\bm{x}}_{1},{\bm{x}}_{2},{\bm{x}}_{3}\right)\,e^{-i{\bm{p}}_{1}\cdot\left({\bm{x}}_{1}-{\bm{x}}_{1}^{\prime}\right)-i{\bm{p}}_{2}\cdot\left({\bm{x}}_{2}-{\bm{x}}_{2}^{\prime}\right)-i{\bm{p}}_{3}\cdot\left({\bm{x}}_{3}-{\bm{x}}_{3}^{\prime}\right)}\,\\ \ \times W_{\mathrm{A}}\left({\bm{p}}_{1},{\bm{p}}_{2},{\bm{p}}_{3},\frac{{\bm{x}}_{1}+{\bm{x}}_{1}^{\prime}}{2},\frac{{\bm{x}}_{2}+{\bm{x}}_{2}^{\prime}}{2},\frac{{\bm{x}}_{3}+{\bm{x}}_{3}^{\prime}}{2}\right)\,.$ (61)
@@
Appendix B HH description of the three-particle wave functions
@@
Regarding hypertriton, we have computed its wave function using the HH method. Since, this application has not reported before, here we give the details of the procedure. In the following, we will specialize to the case where particles $1$ and $2$ are the nucleons, and particle $3$ the $\Lambda$ hyperon. The Jacobi vectors are then defined as
$\left\{\begin{array}[]{l}\bm{\xi}_{2}=\bm{r}_{2}-\bm{r}_{1}\\ \\ \bm{\xi}_{1}=\sqrt{\frac{4\kappa}{2+\kappa}}\ (\bm{r}_{3}-\frac{\bm{r}_{1}+\bm{r}_{2}}{2})\;,\end{array}\right.$ (62)
where $\bm{r}_{i}$ is the position of particle $i$, $\kappa=m_{\Lambda}/M$, $M$ being the nucleon mass and $m_{\Lambda}$ the mass of the $\Lambda$ hyperon (in this paper, we disregard the neutron-proton mass difference). The Jacobi coordinates are combined to construct the hyperspherical coordinates, which contain one radial coordinate, the hyperradius $\rho=(\xi_{1}^{2}+\xi_{2}^{2})^{1/2}$, and five hyperangles (the four angles describing the directions of $\bm{\xi}_{1}$ and $\bm{\xi}_{2}$ plus $\varphi=\arctan(\xi_{1}/\xi_{2})$), which collectively are denoted as $\Omega$.
With these definitions, we have, in the center-of-mass system
$\frac{\rho^{2}}{2}=r_{1}^{2}+r_{2}^{2}+\kappa r_{3}^{2}=\frac{2\kappa}{2+\kappa}(r_{31}^{2}+r_{32}^{2}+\frac{1}{\kappa}r_{12}^{2})$ (63)
with $r_{ij}=|\bm{r}_{i}-\bm{r}_{j}|$ the relative distance between particles $i$ and $j$.
The kinetic energy operator can be written in terms of the hyperspherical coordinates as [37]
$T$ $=$ $-{\hbar^{2}\over 2m}\nabla_{1}^{2}-{\hbar^{2}\over 2m}\nabla_{2}^{2}-{\hbar^{2}\over 2M}\nabla_{3}^{2}\ ,$ (64)
$=$ $-{\hbar^{2}\over m}\left[{d^{2}\over d\rho^{2}}+{5\over\rho}{d\over d\rho}-{K^{2}(\Omega)\over\rho^{2}}\right]\ ,$
where $K^{2}(\Omega)$ is the so-called hyperangular operator, depending only on the variables $\Omega$. The HH functions are the eigenfunctions of this operator, with eigenvalue
$K^{2}(\Omega){\cal Y}_{[K],L,L_{z}}(\Omega)=K(K+4){\cal Y}_{[K],L,L_{z}}(\Omega)\ .$ (65)
Their explicit expression is [37]
${\cal Y}_{[K],L,L_{z}}(\Omega)=N_{[K]}\Bigl{[}Y_{l_{1}}(\hat{\xi}_{1})Y_{l_{2}}(\hat{\xi}_{2})\Bigr{]}_{L,L_{z}}(\cos\varphi)^{l_{2}}(\sin\varphi)^{l_{1}}P^{l_{1}+{1\over 2},l_{2}+{1\over 2}}_{n_{2}}(\cos 2\varphi)\,$ (66)
where $Y_{l}$ are spherical harmonic functions, $P^{a,b}_{n}$ Jacobi polynomials, $L,L_{z}$ the total orbital angular momentum and its $z$ projection, $N_{[K]}$ a normalization factor, and $[K]\equiv\{l_{1},l_{2},n_{2}\}$. Finally, $K$ is the grand angular momentum quantum number, given by $K=l_{1}+l_{2}+2n_{2}$ ($l_{1},l_{2},n_{2}$ are non-negative integers). The HH functions ${\cal Y}_{[K],L,L_{z}}(\Omega)$ form a complete and orthonormal set in the five-dimensional $\Omega$ space.
To construct the hypertriton wave function, we need also to take into account the spin and isospin degrees of freedom. We introduce the basis states
$\Phi_{\mu}=\Bigl{[}{\cal Y}_{[K],L}(\Omega)[(s_{1}s_{2})_{S_{2}}s_{3}]_{S}\Bigr{]}_{J.J_{z}}[t_{1}t_{2}]_{T,T_{z}}\ ,$ (67)
where $s_{i}$ is the spin state of particle $i$. The spin states are coupled to a total spin $S$, which is coupled to the total orbital angular momentum $L$ to form the total angular momentum $J$ of the state we want to describe. The isospin states $t_{1}$ and $t_{2}$ of the two nucleons are coupled to a total isospin $T,T_{z}$ (we remember here that the isospin of the $\Lambda$ hyperon is zero). Above $\mu\equiv\{l_{1},l_{2},n_{2},L,S_{2},S,T\}$. Finally, the hypetriton wave function is written as
$\Psi_{J,J_{z}}=\sum_{\mu}u_{\mu}(\rho)\Phi_{\mu}\,$ (68)
where $u_{\mu}(\rho)$ are functions of the hyperradius to be determined. In order to assure the antisymmetry with respect the exchange of particles $1$ and $2$ (the two nucleons), the sum over the quantum number $\mu$ is limited to the values
$(-)^{l_{2}+S_{2}+T}=-1\ .$ (69)
The hypertriton is a $np\Lambda$ state of total angular momentum and parity $J^{\pi}=1/2^{+}$ and $T=0$. So we have also the additional conditions
$|L-S|\leq J\leq L+S\ ,\qquad(-)^{l_{1}+l_{2}}=+1\ ,\qquad T=T_{z}=0\ .$ (70)
Moreover, the condition to antisymmetrize the two nucleons reduces to $(-)^{l_{2}+S_{2}}=-1$ which means that for the two-nucleons in $s$-wave only the spin $S=1$ is allowed.
The hyperradial functions $u_{\mu}(\rho)$ have been expanded in terms of the following functions [37]
$u_{\mu}(\rho)=\sum_{n=1}^{N}a_{\mu,n}g_{n}(\rho)\ ,\qquad g_{n}(\rho)=N_{n}L^{(5)}_{n-1}(\gamma\rho)e^{-{1\over\rho}\gamma\rho}\ ,$ (71)
Where $N_{n}=\gamma^{3}\sqrt{n!/(n+5)!}$ are normalization factors and $\gamma$ a variational parameter. The coefficients $a_{\mu,n}$ are fixed by means of the Rayleigh-Ritz variational principle, which reads
$\delta{\langle\Psi_{J,J_{z}}|H|\Psi_{J,J_{z}}\rangle\over\langle\Psi_{J,J_{z}}|\Psi_{J,J_{z}}\rangle}=0\ ,$ (72)
where $H$ is the Hamiltonian of the system, which we take of the form
$H=T+V_{NN}(1,2)+V_{N\Lambda}(1,3)+V_{N\Lambda}(2,3)+W_{NN\Lambda}(1,2,3)\ .$ (73)
Above $T$ is the kinetic energy operator given in Eq. (64), while $V_{NN}$ and $V_{N\Lambda}$ are $NN$ and $N\Lambda$ potentials. Finally, $W_{NN\Lambda}$ is a three-particle potential.
In this paper, we have adopted two versions for these interactions. First of all, we have used a number of simple central interactions of the Gaussian form [43]
$V_{NN}(1,2)$ $=$ $\sum_{S=0,1}\overline{V}_{S}e^{-(r_{12}/\overline{r}_{s})^{2}}{\cal P}_{0,S}(1,2)\ ,$ (74)
$V_{N\Lambda}(1,3)$ $=$ $\sum_{S=0,1}V_{S}e^{-(r_{13}/r_{s})^{2}}{\cal P}_{0,S}(1,3)\ ,$ (75)
$W_{NN\Lambda}(1,2,3)$ $=$ $W_{3}\,e^{-(r_{13}^{2}+r_{23}^{2})/\rho_{3}^{2}}\ ,$ (76)
where ${\cal P}_{0,S}(i,j)$ are projectors on the orbital angular momentum $\ell=0$ state and on the spin $S$ state of particles $i$ and $j$. Clearly $V_{N\Lambda}(2.3)$ can be obtained from $V_{N\Lambda}(1,3)$ after the substitution $1\leftrightarrow 2$, so in the present study we do not consider charge-symmetry breaking terms in the $N\Lambda$ interaction. The low-energy Gaussian representation of the $N\Lambda$ interaction is justified in the context of a leading order effective interaction based on contact interactions, see for example Ref. [60].
The parameters $\overline{V}_{S}$ and $\overline{r}_{s}$ have been chosen in order to reproduce the experimental $np$ $s$-wave scattering lengths and effective ranges. The adopted values are: $\overline{V}_{0}=-30.5459$ MeV, $\overline{r}_{0}=1.831$ fm, $\overline{V}_{1}=-66.5825$ MeV, and $\overline{r}_{1}=1.558$ fm. Regarding the $N\Lambda$ systems, such experimental values are not well known, so the parameters $V_{S}$ and $r_{s}$ have been fixed to reproduce the $s$-wave scattering lengths and effective ranges as calculated by the chiral EFT potentials developed in Refs. [47, 48, 49]. In those references different values of the cutoff have been used to regularize the interaction giving in each case different values of the scattering parameters. Accordingly, for each of them we have extracted different values of $V_{S}$ and $r_{s}$ [43]. Furthermore, for each of this Gaussian $V_{N\Lambda}$ we need to include a three-body force, otherwise the hypertriton would be overbound. The corresponding three body parameters $W_{3}$ and $\rho_{3}$ are then chosen to reproduce the hypertriton binding energy $B(^{3}_{\Lambda}{\rm H})=2.40$ MeV [43]. The values of the parameters used in this work are reported in Table 2
Model 1 9
$V_{0}\,$(MeV) -35.130 -31.851
$r_{0}\,$(fm) 1.375 1.434
$V_{1}\,$(MeV) -23.239 -42.055
$r_{1}\,$(fm) 1.482 1.126
$W_{3}\,$(MeV) 12.834 11.795
$\rho_{3}\,$(fm) 2.0 2.0
Table 2: The strength $V_{S}$ and range $r_{S}$ in two spin channels $S=0,1$ of the two Gaussian potentials adopted in this work. These parameters reproduces the $s$-wave scattering lengths and effective ranges as calculated by the chiral EFT potentials devised in The last two rows give the strength and range of the three-body force needed to describe the experimental value of $B(^{3}_{\Lambda}{\rm H})$. The model numbering is taken from Ref. [43].
In addition to the Gaussian interaction, we consider as well the Bodmer-Usmani $N\Lambda$ interaction [51, 52]
$V_{N\Lambda}^{BU}(1,3)=V_{C}(r_{13})(1-\epsilon+\epsilon P_{x})+0.25\,V_{G}\,T^{2}_{\pi}(r_{13})\,\mathbf{\sigma}_{1}\cdot\mathbf{\sigma}_{3},\quad T_{\pi}(r)=\left(1+{3\over\mu r}{3\over(\mu r)^{2}}\right){e^{-\mu r}\over\mu r}\left(1-e^{-c_{\pi}r^{2}}\right)^{3}$ (77)
with $P_{x}$ the space-exchange operator. The $r$-dependence of the potentials has the form $V_{C}(r)=W_{C}/(1+e^{(r-\bar{r})/a})-\bar{v}\,T^{2}_{\pi}(r)$ where $T_{\pi}(r)$ is the pion exchange tensor potential. The values of the parameters used in the present study are: $\epsilon=0$, $\bar{r}=0.5$ fm, $a=0.2$ fm, $\bar{v}=6.20$ MeV, $V_{G}=0.25$ MeV, $W_{C}=2137$ MeV, $\mu=0.69953$ fm-1, and $c_{\pi}=2.1$ fm-2. Associated with this potential we have used the following three-body force
$W_{NN\Lambda}^{BU}(1,2,3)=W_{3}^{BU}T_{\pi}^{2}(r_{13})T_{\pi}^{2}(r_{23})$ (78)
with the value $W_{3}^{BU}=0.025\,$MeV in order to reproduce the hypertrition binding energy.
Once the interaction is chosen, standard numerical methods allow for the evaluation of the matrix elements $H_{\mu,n,\mu^{\prime},n^{\prime}}=\langle g_{n}(\rho)\Phi_{\mu}|H|g_{n^{\prime}}(\rho)\Phi_{\mu^{\prime}}\rangle$, while $\langle g_{n}(\rho)\Phi_{\mu}|g_{n^{\prime}}(\rho)\Phi_{\mu^{\prime}}\rangle=\delta_{n,n^{\prime}}\delta_{\mu,\mu^{\prime}}$ since $\Phi_{\mu}$ and $g_{n}(\rho)$ are orthonomal sets. The application of the Rayleigh-Ritz variational principle therefore reduces to a standard eigenvalue problem
$\sum_{\mu^{\prime},n^{\prime}}H_{\mu,n,\mu^{\prime},n^{\prime}}a_{\mu^{\prime},n^{\prime}}=Ea_{\mu,n}\ .$ (79)
Note that for a central Gaussian potential, it is sufficient to include in the expansion basis only states $\Phi_{\mu}$ with $L=0$ and $S=1/2$. In Table 3, we show examples of convergence. In the considered cases, we have included in the expansion basis all the states $\Phi_{\mu}$ with quantum numbers $L=0$ and $S=1/2$ and $l_{1}+l_{2}+n_{2}\leq K$ (and compatible with the conditions given in Eqs. (6970)). Then, the convergence is studied by increasing the values of $K$ and $N$, the latter being the number of expansion states of the hyperradial functions.
$K$ $N$ $N_{HH}$ Model 1 Model 9 BU
$20$ 24 66 $-2.267$ $-2.268$ $-2.221$
$30$ 24 136 $-2.340$ $-2.340$ $-2.351$
$40$ 24 231 $-2.366$ $-2.366$ $-2.398$
$50$ 24 351 $-2.376$ $-2.376$ $-2.402$
$60$ 24 496 $-2.380$ $-2.380$ $-2.407$
$60$ 28 496 $-2.380$ $-2.380$ $-2.408$
Table 3: Convergence of the hypertriton energy as function of the maximum $K$ of HH functions $\Phi_{\mu}$ and number of functions $g_{n}$ used to expand the hyperradial functions for the various interactions used in this work. Above $N_{HH}$ is the total number of HH states included in the expansion basis for each $K$ value. Here we have performed the calculation for $\gamma=2.8$ fm-1. In all the cases, the Hamiltonian includes the $NN\Lambda$ interaction.
As it can be seen by inspecting Table 3, the convergence is well achieved for values of $K$ around $50$-$60$. Values of the parameter $N$ around $24$-$28$ are sufficient to have an accurate description of the hyperradial functions $u_{\mu}(\rho)$, see Eq. (71). In conclusion, the HH method fournishes accurate description of the hypertriton for the adopted potential models.
Appendix C Congleton Formalism for Hypertriton
For the Congleton formalism, we use the Jacobi coordinate transformation given, and the total position space wave function is given as where $\Psi_{{{}^{3}_{\Lambda}{\rm H}}}$ and $\Psi_{i}$ for $i=1,2,3$ are the wave functions for ${{}^{3}_{\Lambda}{\rm H}}$ and nucleons. $\Psi_{{{}^{3}_{\Lambda}{\rm H}}}$ is assumed to be factorized into the plane wave for COM motion and an internal wave function that depends only on two relative coordinates.
$\Psi_{{{}^{3}_{\Lambda}{\rm H}}}\left(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3}\right)=\left(2\pi\right)^{-3/2}e^{i\mathbf{P}_{{{}^{3}_{\Lambda}{\rm H}}}\cdot\mathbf{R}}\varphi_{{{}^{3}_{\Lambda}{\rm H}}}(\bm{\xi}_{1},\bm{\xi}_{2})$ (80)
Where relative Jacobi coordinates are defined in terms of the mass of two identical masses $m_{1}=m_{2}=M$ for proton and neutron and a third particle, $\Lambda$ with a different mass $m_{3}$. We define the mass ratio as $\kappa=\frac{m_{3}}{M}$
$J^{-1}=\left(\begin{array}[]{ccc}0&\frac{1}{2}&-\frac{1}{2}\\ \frac{2}{2+\kappa}&-\frac{\kappa}{2+\kappa}&-\frac{\kappa}{2+\kappa}\\ \frac{\kappa}{2+\kappa}&\frac{\kappa}{2+\kappa}&\frac{2}{2+\kappa}\\ \end{array}\right)$ (81)
$\begin{pmatrix}\bm{\xi}_{1}\\ \bm{\xi}_{2}\\ \mathbf{R}\end{pmatrix}=J\times\begin{pmatrix}\mathbf{x}_{1}\\ \mathbf{x}_{2}\\ \mathbf{x}_{3}\end{pmatrix}$ (82)
with $|J|=-\frac{(2+\kappa)^{2}}{2+\kappa^{2}}$
$\mathbf{x}_{1}$ $=\frac{\kappa\mathbf{R}}{2+\kappa}+\frac{2\bm{\xi}_{2}}{2+\kappa},\,\,\,\mathbf{x}_{1}^{\prime}=\frac{\kappa\mathbf{R}^{\prime}}{2+\kappa}+\frac{2\bm{\xi}_{2}^{\prime}}{2+\kappa},$ (83)
$\mathbf{x}_{2}$ $=\frac{\mathbf{R}}{2+\kappa}+\frac{\bm{\xi}_{1}}{2}-\frac{\kappa\bm{\xi}_{2}}{2+\kappa},\,\,\,\mathbf{x}_{2}^{\prime}=\frac{\mathbf{R}^{\prime}}{2+\kappa}+\frac{\bm{\xi}_{1}^{\prime}}{2}-\frac{\kappa\bm{\xi}_{2}^{\prime}}{2+\kappa},$
$\mathbf{x}_{3}$ $=\frac{\mathbf{R}}{2+\kappa}-\frac{\bm{\xi}_{1}}{2}-\frac{\kappa\bm{\xi}_{2}}{2+\kappa},\,\,\,\mathbf{x}_{3}^{\prime}=\frac{\mathbf{R}^{\prime}}{2+\kappa}-\frac{\bm{\xi}_{1}^{\prime}}{2}-\frac{\kappa\bm{\xi}_{2}^{\prime}}{2+\kappa},$

### v2

@@ Appendix D Minimum bias multiplicity
$\gamma\frac{\mathrm{d}N_{\mathrm{A}}}{\mathrm{\penalty 10000\ d}^{3}P}=\frac{S_{\mathrm{A}}}{(2\pi)^{14}}|J^{-1}|^{3}\int\mathrm{d}^{3}\bm{q}_{1}\int\mathrm{d}^{3}\bm{q}_{2}\int\mathrm{d}^{3}\bm{S}\int\mathrm{d}^{3}\bm{\xi}_{1}\,\int\mathrm{d}^{3}\bm{\xi}_{2}\int\mathrm{d}^{3}\bm{\xi}_{1}^{\prime}\int\mathrm{d}^{3}\bm{\xi}_{2}^{\prime}\\ \times\varphi^{*}_{NR}\left(\bm{\xi}_{1}^{\prime},\bm{\xi}_{2}^{\prime}\right)\varphi_{NR}\left(\bm{\xi}_{1},\bm{\xi}_{2}\right)\,e^{-i{\bm{q}}_{1}\cdot\left(\bm{\xi}_{1}-\bm{\xi}_{1}^{\prime}\right)-i{\bm{q}}_{2}\cdot\left(\bm{\xi}_{2}-\bm{\xi}_{2}^{\prime}\right)}\,\\ \times W_{\mathrm{A}}\left((c_{1}P^{0}_{d},{\bm{p}}_{1}),(c_{2}P^{0}_{d},{\bm{p}}_{2}),(c_{3}P^{0}_{d},{\bm{p}}_{3}),\frac{(t_{0},{\bm{x}}_{1}+{\bm{x}}_{1}^{\prime})}{2},\frac{(t_{0},{\bm{x}}_{2}+{\bm{x}}_{2}^{\prime})}{2},\frac{(t_{0},{\bm{x}}_{3}+{\bm{x}}_{3}^{\prime})}{2}\right)\Bigr|_{P=P_{\mathrm{A}}}\,.$ (62)
At this point, $t_{0}$ can be taken arbitrarily as $t_{0}=0$. Reinserting the integration over $d^{3}S\times\delta^{(3)}({\bm{R}}-{\bm{R}}^{\prime})$, and going back to the integration over the particle coordinates and momenta, we have
$\gamma\frac{\mathrm{d}N_{\mathrm{A}}}{\mathrm{d}^{3}P}=\frac{S_{\mathrm{A}}}{(2\pi)^{14}}\int\mathrm{d}^{3}\bm{p}_{1}\int\mathrm{d}^{3}\bm{p}_{2}\int\mathrm{d}^{3}\bm{p}_{3}\int\mathrm{d}^{3}{\bm{x}}_{1}\int\mathrm{d}^{3}{\bm{x}}_{2}\,\int\mathrm{d}^{3}{\bm{x}}_{3}\int\mathrm{d}^{3}{\bm{x}}_{1}^{\prime}\int\mathrm{d}^{3}{\bm{x}}_{2}^{\prime}\int\mathrm{d}^{3}{\bm{x}}_{3}^{\prime}\\ \times\Psi_{NR}^{*}\left({\bm{x}}_{1}^{\prime},{\bm{x}}_{2}^{\prime},{\bm{x}}_{3}^{\prime},\right)\Psi_{NR}\left({\bm{x}}_{1},{\bm{x}}_{2},{\bm{x}}_{3}\right)\,e^{-i{\bm{p}}_{1}\cdot\left({\bm{x}}_{1}-{\bm{x}}_{1}^{\prime}\right)-i{\bm{p}}_{2}\cdot\left({\bm{x}}_{2}-{\bm{x}}_{2}^{\prime}\right)-i{\bm{p}}_{3}\cdot\left({\bm{x}}_{3}-{\bm{x}}_{3}^{\prime}\right)}\,\\ \ \times W_{\mathrm{A}}\left({\bm{p}}_{1},{\bm{p}}_{2},{\bm{p}}_{3},\frac{{\bm{x}}_{1}+{\bm{x}}_{1}^{\prime}}{2},\frac{{\bm{x}}_{2}+{\bm{x}}_{2}^{\prime}}{2},\frac{{\bm{x}}_{3}+{\bm{x}}_{3}^{\prime}}{2}\right)\,.$ (63)
@@
Appendix F HH description of the three-particle wave functions
@@
Appendix G Congleton Formalism for Hypertriton
For the Congleton formalism, we use the Jacobi coordinate transformation given in Ref. [34], and the total position space wave function is given as where $\Psi_{{{}^{3}_{\Lambda}{\rm H}}}$ and $\Psi_{i}$ for $i=1,2,3$ are the wave functions for $^{3}_{\Lambda}{\rm H}$} and nucleons. $\Psi_{{{}^{3}_{\Lambda}{\rm H}}}$ is assumed to be factorized into the plane wave for center-of-mass motion and an internal wave function that depends only on two relative coordinates.
$\Psi_{{{}^{3}_{\Lambda}{\rm H}}}\left(\bm{x}_{1},\bm{x}_{2},\bm{x}_{3}\right)=\left(2\pi\right)^{-3/2}e^{i\bm{P}_{{{}^{3}_{\Lambda}{\rm H}}}\cdot\bm{R}}\varphi_{{{}^{3}_{\Lambda}{\rm H}}}(\bm{\xi}_{1},\bm{\xi}_{2})\ ,$ (64)
where the relative Jacobi coordinates are defined in terms of the mass of two identical particles $m_{1}=m_{2}=M$ for proton and neutron and a third particle, $\Lambda$ with a different mass $m_{3}$. We define the mass ratio as $\kappa=\frac{m_{3}}{M}$ and writing single particle position and momentum cordinates $\bm{x}_{i}$, $\bm{p}_{i}$ for $i=1,2,3$
$\bm{x}_{1}$ $=\frac{\kappa\bm{R}}{2+\kappa}+\frac{2\bm{\xi}_{2}}{2+\kappa},\,\,\,\bm{x}_{1}^{\prime}=\frac{\kappa\bm{R}^{\prime}}{2+\kappa}+\frac{2\bm{\xi}_{2}^{\prime}}{2+\kappa},$ (65)
$\bm{x}_{2}$ $=\frac{\bm{R}}{2+\kappa}+\frac{\bm{\xi}_{1}}{2}-\frac{\kappa\bm{\xi}_{2}}{2+\kappa},\,\,\,\bm{x}_{2}^{\prime}=\frac{\bm{R}^{\prime}}{2+\kappa}+\frac{\bm{\xi}_{1}^{\prime}}{2}-\frac{\kappa\bm{\xi}_{2}^{\prime}}{2+\kappa},$
$\bm{x}_{3}$ $=\frac{\bm{R}}{2+\kappa}-\frac{\bm{\xi}_{1}}{2}-\frac{\kappa\bm{\xi}_{2}}{2+\kappa},\,\,\,\bm{x}_{3}^{\prime}=\frac{\bm{R}^{\prime}}{2+\kappa}-\frac{\bm{\xi}_{1}^{\prime}}{2}-\frac{\kappa\bm{\xi}_{2}^{\prime}}{2+\kappa},$